OpenStudy (anonymous):

find the equation, in standard form, of the line passing through the points (5,1) and (-2,3)

OpenStudy (whpalmer4):

First thing to do is to find the slope of the line passing through those points, call them \((x_1,y_1)\) and \((x_2,y_2)\). Slope (\(m\)) is given by \[m =\frac{y_2-y_1}{x_2-x_1}\] What do you get?

OpenStudy (anonymous):

2/-7

OpenStudy (whpalmer4):

Okay, that sounds right. Next, take either known point, and use the slope you just found in the point-slope formula: \[y-y_0 = m(x-x_0)\]That will give you an equation for the line passing through the point \((x_0,y_0)\) with slope \(m\). Finally, rearrange the resulting equation into standard form, \(Ax+By=C\), with \(A>0\) and \(A, B, C\) relatively prime integers. In other words, you shouldn't be able to factor out any common factors from the equation.